### 11 Jan 2024 v2

# Optimal map-making with singularities

Abstract
In this work, we investigate the optimal map-making technique for the linear system $\bm{d}=\bm{A}\bm{x}+\bm{n}$ while carefully taking into account singularities that may come from either the covariance matrix $\bm{C}=\langle\bm{n}\bm{n}^{t}\rangle$ or the main matrix $\bm{A}$. We first describe the general optimal solution, which is quite complex, and then use the modified pseudo inverse to create a near-optimal solution, which is simple, robust, and can significantly alleviate the unwanted noise amplification during map-making. The effectiveness of the nearly optimal solution is then compared to that of the naive co-adding solution and the standard pseudo inverse solution, showing noticeable improvements. Interestingly, all one needs to get the near-optimal solution with singularity is just a tiny change to the classical solution, which is designed for the case without singularity.
1 Introduction
The precise measurement of the temperature and polarization of the Cosmic Microwave Background (CMB) has allowed us to establish the standard cosmological model ($\Lambda$CDM). The theory of inflation was developed in the 1980s to address the flatness, horizon, and monopole concerns within the $\Lambda$CDM framework [1, 2]. Despite inflationary cosmology’s enormous success, the expected primordial gravitational waves (PGW) have not been discovered yet. PGW detection and inflationary cosmology testing remain critical parts of current and future CMB researches. The most promising technique to detect the PGW is via the CMB B-mode polarization observations, according to [3, 4, 5].
With the completion of the Planck mission [6], the fundamental scientific goals of various present and future CMB studies are set to detecting the CMB B-mode polarization. These include the ground-based experiments like BICEP/Keck [7, 8], SPTpol [9], CLASS[10], ACTpol [11], POLARBEAR [12], and AliCPT [13, 14, 15] in the Northern Hemisphere; as well as the balloon experiments like EBEX [16], SPIDER [17] and future experiments like SO [18], CMB-S4 [19], and the space mission LiteBIRD [20], which is scheduled to be launched in 2029.
To obtain a more useful CMB polarization dateset, the time-ordered data (TOD) is often compressed into sky maps, and the data volume is reduced from tens of trillions to tens of millions, which is called a mapmaking approach. This kind of approach seeks to condense the data while maintaining as much cosmological information as possible. Typically, mapmaking is treated as a linear problem with the goal of generating an unbiased estimate of the CMB sky map while minimizing its variance. This is usually accomplished using approaches such as the minimum variance or maximum likelihood estimation, e.g., [21, 22, 23, 24, 25, 26, 27].
However, in the case of real TOD, a number of time-domain filtering processes are frequently required to remove the atmosphere and ground radiation, as well as numerous systematics. Although these operations are necessary, they have an inevitable side effect of removing a considerable percentage of the desirable signal at the same time. As a consequence, some of the signal modes are permanently lost due to the filtering and, if the filtering consists of operations in multiple domains such as Fourier and polynomial, some of the signal modes may be suppressed rather than removed, leading to a reduced signal-to-noise ratio (SNR) and an amplified noise in the final sky map.
For a naive map-making algorithm like co-adding and averaging, the above mentioned issues are not big problems, because it does not pay much attention to further improving the SNR in the final sky map product – which is done by optimized map-making algorithms. Unfortunately, most optimized map-making algorithms require to use the inversion of the covariance matrix, which is problematic if singularities are involved. To the best of our knowledge, there is still no comprehensive discussion regarding a rigorous optimal solution of map-making with the presence of various singularity problems, despite the fact that there are some applicable solutions to this issue with certain compromises, such as using a modest addition of human choice to make the covariance matrix non-singular.
In this work, we will explore a rigorous optimal map-making method that takes into account various singularity concerns. We will also show how to simplify the optimal solution while tightly limiting any unintended side effects along with the simplification. The final recommended solution is simple, near-optimal, and can significantly alleviate the above mentioned noise amplification effect.
The outline of this work is the following: We first introduce the optimal and near-optimal solutions in section 2, with the corresponding lengthy mathematical details put in appendix A; then we discuss some technical details of the solution in section 3, followed by the code validation and simulation tests in section 4. Finally, we give a conclusion and some further discussions in section 5.
2 Methodology
Because the mathematical deduction of the optimal solution of map-making with singularity is quite lengthy, we provide only a brief introduction in this section, and put the main contents in appendix A.
2.1 The classical minimum variance solution and its limitations
The problem starts from the well-known matrix equation that connect the pixel domain sky map $\bm{x}$ to the time-order data $\bm{d}$ (TOD) with noise $\bm{n}$:
$\bm{d}=\bm{A}\bm{x}+\bm{n},$ (2.1)
where $\bm{x}$ is a column vector of size $n_{\rm{pix}}$; $\bm{n}$ is another column vector of size $n_{\rm tod}$; and $\bm{A}$ is a matrix with $n_{\rm tod}$ rows and $n_{\rm{pix}}$ columns. In order to increase the signal to noise ratio (SNR), $n_{\rm tod}$ is usually much bigger than $n_{\rm{pix}}$, so each sky pixel is observed many time. The goal of map-making is to find the best solution of $\bm{x}$ from $\bm{d}$, and the simplest case of the solution requires the following:
  1. 1.
    Each row of $\bm{A}$ contains only one element that is equal to 1, and all other elements are zero. Thus, $\bm{A}^{t}\bm{A}=\bm{N}_{obs}$ is diagonal [^1] Otherwise at least one row will contain more than one elements that are equal to 1 and contains the number of observations at each sky pixel.
  2. 2.
    $\bm{A}$ has full column-rank and $\bm{A}^{t}\bm{A}$ is invertible.
  3. 3.
    The noise covariance matrix $\bm{C}=\langle\bm{n}\bm{n}^{t}\rangle$ is invertible and fully represents the noise property.
Given the fulfillment of these conditions, the optimal solution $\widetilde{\bm{x}}$ for the sky signal is given by the equation below, which has been documented in various publications, e.g., [28, 29]:
$\widetilde{\bm{x}}=(\bm{A}^{t}\bm{C}^{-1}\bm{A})^{-1}\bm{A}^{t}\bm{C}^{-1}\bm{d}.$ (2.2)
Certainly, in instances where the aforementioned conditions are not satisfied, particularly regarding items 2 and 3, the above solution is no longer applicable. Especially, when the singularity exists in either or both of $\bm{A}^{t}\bm{A}$ and $\bm{C}$, the true optimal solution can be complicated, which has not been documented in detail before.
2.2 The general minimum variance solution with singularity
In this work, we give the optimal map-making solution in general, which requires only two fundamental conditions that are almost always true:
  1. 1.
    $n_{\rm tod}\gg n_{\rm{pix}}$, so each sky pixel is observed many times.
  2. 2.
    The noise covariance matrix $\bm{C}=\langle\bm{n}\bm{n}^{t}\rangle$ does converge.
Apparently, such a significant relaxation of the conditions means $\bm{A}^{t}\bm{A}$ and $\bm{C}$ are both allowed to be singular, and no constraint is assumed for the relationship between $\bm{A}$ and $\bm{C}$. A lengthy introduction of how to get the general optimal solution under the aforementioned conditions can be found in appendix A. In summary, the mathematical inference yields two key findings:
First of all, it is possible to obtain a general true optimal solution, which requires to use the singular value decomposition to carefully analyze the linear system and the origin of singularity, in order to clearly separate the singularity that arise from various origins, and use all available information properly to obtain the final optimal solution. Details of the true optimal solution can be found in appendix A.1 – A.3, and the main process includes first dealing with the covariance matrix’s singularity, and then treating the main matrix’s singularity properly, to obtain the final optimal solution by eq. (A.101).
Secondly, since the true optimal solution is complicated and uneasy to use, we have designed a carefully simplified solution based on the modified pseudo inverse, as introduced in appendix A.5.2 and particularly, in eq (A.119). The simplified solution requires only a tiny change to the classical pseudo inverse solution, but has three major advantages: A) It is much easier to use, because the solution is contained in a single equation. B) The side effect due to simplification is strictly limited to the singular part, which is usually negligible (see appendix A.6 for more discussions). C) The solution can significantly alleviate the unwanted noise amplifications (see section 3.1 for more details).
3 A few discussions about the technical details
In this section, we discuss a few technical details, which do not affect the theoretical aspects of the solution, but may be useful for pursuing the best effect in an implementation of the method.
3.1 The noise amplification effect in optimal mapmaking
Another strong reason to adopt the near optimal solution in eq. (A.119) is to prevent the noise amplification effect in optimal mapmaking, which is tightly associated with the overall scan strategy (structure of $\bm{A}$) and the properties of filtering. Below we discuss this phenomenon and explain why the near optimal solution helps enormously to prevent such an unwanted effect.
In the case where matrix $\bm{A}$ possesses full column rank, the matrix $\bm{A}^{t}\bm{C}^{\times}\bm{A}$ always remains non-singular and is safe to use; however, the behavior of matrix $\bm{A}^{t}\bm{C}^{+}\bm{A}$ (compute the inversion of $\bm{C}$ with the standard pseudo inverse) is impacted by the structure of matrix $\bm{A}$. For example, consider a special case where celestial observations are undertaken in a peculiar manner such that all pixels sharing the same signal value are put into one group, and the process of filtering consists merely of excluding the average value of each group. Apparently, this kind of strategy will remove the signal completely, whereas noise continues to prevail, leading to a zero SNR for all pixel domain modes after filtering.
Certainly, the above example will not happen in reality, but it does tell us that, the joint effect of scan strategy and TOD filtering can possibly cause a deteriorated SNR. Meanwhile, the minimum variance solution will try to maintain signal integrity as far as possible. Therefore, if the filtering causes a significant SNR deterioration at a particular pixel domain mode (provided it doesn’t plummet to zero), the minimum variance solution will automatically amplify the total signal to keep the desired component lossless, which leads to an inevitable noise amplification in this mode.
The above mentioned noise amplification effect is characterized by the small eigenvalues of $\bm{A}^{t}\bm{C}^{+}\bm{A}$ (not the eigenvalues of $\bm{C}$). Therefore, during the computation of $\bm{A}^{t}\bm{C}^{+}\bm{A}$’s pseudo-inverse, it becomes essential to choose an appropriate cutoff of the small eigenvalues to exclude the components with poor SNR after filtering. An alternative way to solve this problem is to apply a posterior Wiener filter to optimally assign weights to the components according to their SNR. However, as a well known effect, the Wiener filtered signal is no longer lossless.
On the contrary, although the same problem may also exist in the inversion of $\bm{A}^{t}\bm{C}^{\times}\bm{A}$ (the solution with modified pseudo inverse), the chance to get a very small eigenvalue of $\bm{A}^{t}\bm{C}^{\times}\bm{A}$ is greatly reduced, because we have eliminated all zero modes of $\bm{C}$ via the modified pseudo inverse, which make the solution with modified pseudo inverse much more robust.
On further preventing the noise amplification effect, the pointing matrix $\bm{A}$ should, in the ideal case, be composed of random observations. This randomness ensures that the power loss caused by filtering is dispersed evenly across different pixel domain modes, preventing any mode from experiencing a severe SNR decrease, and hence eliminates the aforementioned problem. Although implementing such a random scanning scheme is impractical, it is still helpful to add multiple modes to the scan strategy to alleviate the noise amplification effect, and to make the strategy more robust. Meanwhile, it is crucial to take $\bm{A}^{t}\bm{C}^{+}\bm{A}$’s eigenvalue threshold into account whenever one needs to deal with the $\bm{S}_{+}$ solution.
3.2 Approximation of the noise covariance matrix
In order to achieve the optimal solution, it is desirable to possess knowledge of the noise covariance matrix. Nonetheless, acquiring the true noise covariance matrix presents considerable challenges, particularly in the context of ground-based experiments, where the noise is significantly influenced by atmospheric conditions that vary over time. Consequently, it becomes necessary to consider an approximation of the noise covariance matrix.
Indeed, the simplest approach of map-making involves estimating the signal of a single sky pixel by taking the average of all observations corresponding to that pixel. This method in fact assumes a noise covariance matrix with identical diagonal elements, representing a rather basic and unsophisticated approximation. A more refined technique involves estimating the Fourier spectrum of the noise over a specific time period, thereby incorporating the two-point covariance of the noise. Utilizing a longer time period is generally advantageous as it allows for a greater number of samples of the two-point covariance and a wider range of correlation lengths. In essence, the improvement scales approximately as the square root of the number of points, denoted as $N_{p}$, used in the time segment to estimate the Fourier spectrum. Thus, the aforementioned unsophisticated approximation corresponds to setting $N_{p}$ equal to 1. By considering $N_{p}$ to be the number of TOD points within a given time interval (e.g., a few minutes or thousands of points), we already achieve a substantial enhancement in the estimation of the noise covariance matrix. Further improvements obtained by considering hours of TOD are likely to yield only a marginal enhancement to the estimation of the noise covariance matrix.
In a more comprehensive approach, it is advisable to incorporate the atmospheric emission model while also accounting for the temporal variability of atmospheric emissions and the spatial distribution of such emissions in the local zenith coordinates. However, it should be noted that these considerations extend beyond the scope of this work.
A fundamental aspect to consider when approximating the noise covariance matrix is the requirement that the resultant estimation matrix $\bm{M}$, leading to $\widetilde{\bm{x}}=\bm{M}\bm{d}$, should not possess excessively large singular values. The presence of such large singular values can lead to substantial errors when they happen to interact with the uncertainty inherent in the noise covariance matrix.
4 Code validation, simulation and tests
In ground-based CMB experiments, the methods for converting TOD into maps primarily involve the naive map-making, as used by BICEP [30, 31], POLARBEAR [32], and SPT [33]); and also the maximum likelihood map-making, as employed by ACT [34, 35] and POLARBEAR [26]). While naive map-making is prone to producing imperfect estimates of the sky signal, it can be improved by fine-tuning the filters to match the data property, which enables a more diagonal time-domain noise covariance matrix after filtering. Next, signal-only simulations can be used to correct the angular power spectrum suppression cause by filtering [30, 32, 33]. On the other hand, maximum likelihood map-making, particularly as used by POLARBEAR [26], produces maps with a better signal-to-noise ratio by simultaneously solving for the intended signal and contamination templates, and ACT processes the filtered TOD similarly to unfiltered data [34], treating filters as having little effect on the sky signal.
Both the naive and maximum likelihood map-making algorithms will be tested in this section, and their performance will be compared to the recommended modified pseudo inverse solution, which is introduced in section 2 and detailed in appendix A.5.2. For convenience, we adopt the following symbols: the maximum likelihood map-making (implemented by the pseudo inverse) is indicated as $\bm{S}_{+}$, the modified pseudo inverse solution is represented as $\bm{S}_{\times}$, and the simplest naive map-making is denoted as $\bm{S}_{c}$. The following can be used to express these three solutions:
$\bm{S}_{\rm c}$ $=(\bm{A}^{t}\bm{A})^{-1}\bm{A}^{t}\bm{d}$ (4.1)
$\bm{S}_{+}$ $=(\bm{A}^{t}\bm{C}^{+}\bm{A})^{+}\bm{A}^{t}\bm{C}^{+}\bm{d}$
$\bm{S}_{\times}$ $=(\bm{A}^{t}\bm{C}^{\times}\bm{A})^{+}\bm{A}^{t}\bm{C}^{\times}\bm{d}$
Here, the matrix $\bm{A}$ is determined by the scanning strategies, and we assume that $\bm{A}^{t}\bm{A}$ is non-singular, as is typical in most experiments. Our simulation is based on the observation strategies proposed by AliCPT [13], which include two scanning strategies: a large-area scan for the Milky Way observation, and a small-area scan for the CMB observation. Both scanning strategies made use of the Constant Elevation Scan (CES) mode to guarantee a stable atmospheric payload on detectors.
A code validation was done prior to the simulation, as introduced in Appendix D. Then, sections 4.1 and 4.2 below presents the simulation tests with the above two scan strategies.
4.1 The performance tests: for a large sky region
For the large-area scan simulation, the noise covariance matrix is assumed to be diagonal in the Fourier domain, and its diagonal components are represented by the noise’s Fourier spectrum. Thus, the time domain noise covariance matrix is $\bm{C}=\bm{W}\bm{\lambda}^{2}\bm{W}^{H}$, where $\bm{W}$ and $\bm{W}^{H}$ signify the Fourier transformation matrix and its conjugate transpose respectively, and each column of $\bm{W}$ representing one Fourier mode. The simulated TOD includes the noise computed from $\bm{C}$ and the CMB signal constructed from the Planck 2018 best-fit cosmological parameters [36]. According to the Kolmogorov model of turbulence, the power-law index of the integrated atmospheric emission (noise) varies between $-11/3$ and $-8/3$ [37]; thus, its Fourier spectrum is modeled as
$P(\nu)=\sigma_{0}^{2}\left(1+\left(\frac{f}{f_{\rm knee}}\right)^{\alpha}\right),$ (4.2)
wherein $\nu$ represents the frequency while $\sigma_{0}^{2}$ is the variance of white noise, and $f_{\rm{knee}}$ is the knee frequency. According to the Kolmogorov model of turbulence, the power-law index of the integrated atmospheric emission (noise) varies between $-11/3$ and $-8/3$, thus here we set $\alpha=-3$. For other parameters, we considered an equivalent detector that the noise has been averaged by multi-detectors. Thus, we set the amplitude of white noise $\sigma_{0}^{2}=10^{2}~{}\mu{\rm K}^{2}~{}{\rm Hz}^{-1}$ while the knee frequency $f_{\rm knee}=2.15~{}{\rm Hz}$. When the frequency is less than 0.01 Hz, $P(\nu)$ is constant, which aims to avoid the singularity at $\nu=0$. The power spectrum density for both CMB and atmospheric noise are shown in Figure 1. As elucidated in Section 3.2, the atmospheric emission simulation employed in this study neglects the spatial correlation among the scan rings, leading to a simplification of the noise covariance. It is crucial to emphasize that the noise model we have utilized solely takes into account temporal correlations, while disregarding spatial correlations, as previously discussed in section 3.2.
Figure 1: The power spectrum density for both CMB and atmospheric noise. The vertical doted line shows the threshold frequency. For the singular covariance, the power spectrum is set to be 0 at frequencies less than the threshold.
The large-area scan utilizes a basic circular scanning strategy with a fixed elevation of $45^{\circ}$. The TOD sampling frequency is fixed at 50 Hz with the circular scan enduring for a period of one minute. Under these conditions, and with a reconstructed map resolution of $N_{side}=128$, we obtain a total of $n_{\rm{pix}}=706$ pixels in the pixel domain, covering approximately $f_{\rm sky}=35\%$ of the sky. To accomplish this, the 8-hour observation is divided into 19 scan-sets, each spanning a duration of $25$ minutes. By assuming the absence of correlation between these scan-sets, independent solutions are derived from each one. Eventually, a co-addition is performed on these scan-sets, employing an inverse noise variance weight in the pixel domain. Neglecting the variance of the Cosmic Microwave Background (CMB), the inverse noise variance weights are determined as follows:
$w_{ij,\rm c}$ $=\frac{[(\bm{A}^{t}_{j}\bm{A}_{j})^{-1}]^{-1}_{ii}}{\sum_{j}[(\bm{A}^{t}_{j}\bm{A}_{j})^{-1}]^{-1}_{ii}},\;w_{ij,\times}=\frac{[(\bm{A}^{t}_{j}\bm{C}^{\times}\bm{A}_{j})^{+}]^{-1}_{ii}}{\sum_{j}[(\bm{A}_{j}^{t}\bm{C}^{\times}\bm{A}_{j})^{+}]^{-1}_{ii}},\;w_{ij,+}=\frac{[(\bm{A}^{t}_{j}\bm{C}^{+}\bm{A}_{j})^{+}]^{-1}_{ii}}{\sum_{j}[(\bm{A}^{t}_{j}\bm{C}^{+}\bm{A}_{j})^{+}]^{-1}_{ii}},$ (4.3)
where $i$ is the pixel index, $j$ is the scan-set index, and subscript $ii$ denotes a matrix’s $i$-th diagonal element. The final sky map is the weighted average of the solutions from each scan-set:
$\hat{x}_{i}=\sum_{j}w_{ij}\hat{x}_{ij}$ (4.4)
In the context of a singular covariance matrix $\bm{C}$, the map-level outcomes are depicted in Figure 2, wherein the residual of the $\bm{S}_{\times}$ solution is visually lower compared to the other two solutions.
Figure 2: Upper panel: the maps reconstructed by the solutions $\bm{S}_{\rm c},\bm{S}_{\times}$ and $\bm{S}_{+}$ (from left to right). Lower panel: the corresponding pixel domain residuals (output - noiseless CMB input). Obviously, the result by $\bm{S}_{\times}$ (middle) has the lowest overall residual (lower-middle). The map resolution is $N_{side}=256$.
The angular power spectra of the pixel domain results are derived from an ensemble of 400 simulations and given in Figure 3, which shows that, after eliminating the average noise spectrum and correcting the suppression effect arising from filtering (pertaining exclusively to the $\bm{S}_{+}$ solution) and masking (pertaining to all solutions), unbiased angular power spectra can be obtained from all solutions. Notably, the $\bm{S}_{\times}$ solution displays significantly reduced error bar amplitudes in comparison to both the $\bm{S}_{+}$ solution (for $\ell<80$) and the $\bm{S}_{\rm c}$ solution (for $\ell>30$). Therefore, the $\bm{S}_{\times}$ solution outperforms the other two solutions across the majority of $\ell$-ranges.
Figure 3: The angular power spectrum results corresponding to Figure 2, with the mean noise spectra removed and the suppression factors at each $\ell$ corrected. Top panel: the corrected spectra of one realization as example. Middle panel: difference between the expected CMB angular power spectrum and each solutions’ mean spectrum of 400 simulations. Bottom panel: the corresponding error bar amplitudes at each $\bm{\ell}$.
4.2 The performance tests: for a small sky region
In this validation procedure, we focus on a relatively small sky region, covering approximately $f_{\rm sky}=5\%$. We also used observations with a duration of 8 hours, but they were divided into 16 scan-sets, each lasting 30 minutes. Similarly, we ignore the correlation between scan-sets.
In an actual observational experiments, the scanning strategy is not circular. Instead, the telescope scanned back-and-forth at a fixed elevation angle. In this case, we pay more attention to the spatial correlation of atmospheric emission.
Assuming the atmosphere is a vast 2D plane and the power spectrum of the emission intensity satisfies $P(\bm{k})=\left<f(\bm{k})f^{*}(\bm{k}^{\prime})\right>\propto|\bm{k}|^{-3}\delta(\bm{k}-\bm{k}^{\prime})$, where $f(\bm{k})$ is the 2D Fourier transformation of the atmospheric emission intensity. The covariance in real domain between any two points can be derived trivially:
${\rm Corr}(\delta r)=\left<T(\bm{r})T(\bm{r}+\delta\bm{r})\right>\propto\int_{0}^{\infty}{\rm d}kk^{-2}J_{0}(k\delta r),$ (4.5)
where $J_{0}$ is the first kind Bessel function. The covariance can be obtained through numerical computation. To avoid integral divergence, we also set a cutoff frequency, below which the power spectrum of atmospheric emission is constant. This is similar to what is shown in Figure 1.
The physical distance between two sample points can be easily calculated from scan strategy data. Taking into account the influence of wind on the atmosphere, the covariance on TOD domain can be written as:
$C_{ij}={\rm Corr}(\big{\lvert}\delta\bm{r}_{ij}+\bm{v}_{w}\delta t_{ij}\big{\rvert}),$ (4.6)
where $\delta\bm{r}_{ij},\delta t_{ij}$ denotes the physical distance and time difference between two samples, respectively. $\bm{v}_{w}$ denotes the wind speed. For the sake of simplicity, it is treated as a constant in the simulation.
Any linear filter $\bm{M}$ operating in the TOD domain will also distort the covariance matrix. The relationship between the filtered and unfiltered covariance is simply $\tilde{\bm{C}}=\bm{M}\bm{C}\bm{M}^{t}$. Here, we also employ a high-pass filter with the same parameters as in the previous section. The only difference is that the filter operating on each half-scan independently, which means the filter matrix $\bm{M}$ is a block diagonal matrix.
The filtered covariance $\tilde{\bm{C}}$ is singular due to $\bm{M}$ has some zero eigenvalues. The subsequent map-making processing is similar to that mentioned in the previous section. The map-level results are depicted in Figure 4, along with the corresponding residuals for each solution. The angular power spectra results are illustrated in Figure 5.
Figure 4: The reconstructed map and the corresponding pixel domain residuals. Figure conventions are similar to those of Figure 2. Obviously, the result by $\bm{S}_{\times}$ (middle) has the lowest overall residual (lower-middle). The map resolution is $N_{side}=128$.
Figure 5: The angular power spectrum results corresponding to Figure 4, with the mean noise spectra removed and the suppression factors at each $\ell$ corrected. Top panel: the corrected spectra of one realization as example. Middle panel: difference between the expected CMB angular power spectrum and each solutions’ mean spectrum of 400 simulations. Bottom panel: the corresponding error bar amplitudes at each $\bm{\ell}$.
5 Summary and discussions
In this paper, we show how to solve for the sky map in an optimal way with a full consideration of the singularity problems. In summary, the singular value decomposition, along with a detailed investigation of the effective noise covariance matrix, should be employed to discover the general true optimal solution. On the other hand, a sub-optimal solution that is just slightly different from the classical pseudo inverse method, was designed to greatly simplify the true optimal solution. Simulations with the AliCPT observation strategies confirm that the sub-optimal solution is capable of producing significantly better sky maps and angular power spectra than the naive and classical pseudo inverse solutions, especially at the large scales.
In the list below, we compare the sub-optimal solution $\bm{S}_{\times}$ and the similar solution that uses only the pseudo-inverse, $\bm{S}_{+}$:
  1. 1.
    When $\bm{A}^{t}\bm{A}$ and $\bm{C}$ are both non-singular: Both $\bm{S}_{\times}$ and $\bm{S}_{+}$ are optimal, and the signal is lossless in the solution.
  2. 2.
    When $\bm{A}^{t}\bm{A}$ is singular but $\bm{C}$ is non-singular: $\bm{S}_{\times}$ is slightly sub-optimal, $\bm{S}_{+}$ is optimal. However, none of them can keep the signal lossless.
  3. 3.
    When $\bm{A}^{t}\bm{A}$ is non-singular but $\bm{C}$ is singular: $\bm{S}_{\times}$ is near optimal with strictly limited negative effect, and keeps the signal lossless; $\bm{S}_{+}$ is non- optimal and will hurt the signal. $\bm{S}_{\times}$ is hence almost always better than $\bm{S}_{+}$.
  4. 4.
    When $\bm{A}^{t}\bm{A}$ and $\bm{C}$ are both singular: Can be one of above three.
In all cases, the side effect of $\bm{S}_{\times}$ is strictly limited, and can keeps the signal lossless as much as possible. This is why we tends to adopt $\bm{S}_{\times}$ as the recommended solution.
However, we need to point out the following fact: as mentioned in section 3.1, when the mapmaking algorithm is designed to keep the signal lossless (as a precondition), there is going to be a significant amplification of the noise term in the modes with poor SNR, and the overall mapmaking result may looks bad. This is an inevitable consequence of all “lossless” or “make lossless at all costs” mapmaking algorithms. The solution is also simple: one needs to consider a posterior Wiener filtering style approach to put a reasonable balance between the signal loss and noise control.
In principle, it is even possible to further augment the method in this work, e.g., to solve for the best estimation of the CMB angular power spectrum directly from the TOD, which means to re-write the main equation into
$\bm{d}=\bm{A}\bm{M}\bm{Y}\bm{a}+\bm{n},$ (5.1)
where $\bm{M}$ is a mask corresponding to the observation region, $\bm{Y}$ is a matrix of the spherical harmonics, and $\bm{a}$ is the spherical harmonic coefficients, which satisfy $\langle\bm{a}\bm{a}^{t}\rangle=\bm{C}_{\ell}$ (diagonal), i.e., the covariance matrix of $\bm{a}$ is the CMB angular power spectrum. Then we can either solve for the contribution of $\bm{d}$ to $\bm{a}$ by the method described in this work, or even use an quadratic maximum likelihood (QML) method to solve for the best $\bm{C}_{\ell}$, because the covariance matrix of $d$ is
$\langle\bm{d}\bm{d}^{t}\rangle=\bm{A}\bm{M}\bm{Y}\bm{C}_{\ell}\bm{Y}^{H}\bm{M}\bm{A}^{t}+\langle\bm{n}\bm{n}^{t}\rangle,$ (5.2)
whose partial derivative to $C_{\ell}$ is easy to obtain, and hence enables the QML calculations. Ideally, this can give the best estimation of $C_{\ell}$ even without map-making; however, because we need to work in the pixel domain for the foreground removal and delensing, such a direct solution of $C_{\ell}$ is not a good choice unless those pixel domain operations can be successfully transported to the time domain, which is an exciting idea but does not seems to be feasible in the near future.
The concept of the near-optimal solution can readily assist in the solution of singularity issues in other methods, such as the maximum likelihood solution, so long as the method calls for the inversion of a matrix with a singularity. Simply use the modified pseudo-inverse for the inversion of the noise covariance matrix and the standard pseudo-inverse for all other matrix inversions. The benefits of the near-optimal solution presented in this study are then fully realized.
Acknowledgments
This work is supported by the National Key R&D Program of China No. 2021YFC2203100 and No. 2021YFC2203104, the Anhui Provincial Natural Science Foundation 2308085MA30 and the Anhui project Z010118169.
Appendix A The mathematical details
A.1 How to deal with the noise covariance matrix’s singularity
To deal with a singular noise covariance matrix might be the most complicated part, but in most cases, the solution can be eventually simplified. We shall first introduce the general method that is somewhat complex, and then discuss simplifications to the method.
Because the covariance matrix $\bm{C}=\langle\bm{n}\bm{n}^{t}\rangle$ does converge, it can always be decomposed into $\bm{C}=\bm{P}\bm{\lambda}^{2}\bm{P}^{t}$, where $\bm{P}$ is unitary and $\bm{\lambda}$ is diagonal (allow zero diagonal elements). Therefore, it is always possible to convert $\bm{n}$ into a standard white noise $\bm{n}_{0}$ that follows the standard normal distribution and satisfy $\langle\bm{n}_{0}\bm{n}_{0}^{t}\rangle=\bm{I}$. The connection is given by $\bm{n}=\bm{P}\bm{\lambda}\bm{n}_{0}$, which automatically gives $\bm{C}=\langle\bm{n}\bm{n}^{t}\rangle=\bm{P}\bm{\lambda}^{2}\bm{P}^{t}$; thus, we get
$\bm{d}$ $=\bm{A}\bm{x}+\bm{P}\bm{\lambda}\bm{n}_{0}$ (A.1)
$\bm{P}^{t}\bm{d}$ $=\bm{P}^{t}\bm{A}\bm{x}+\bm{\lambda}\bm{n}_{0}.$
Because $\bm{\lambda}$ may have zero diagonal element, we divide $\bm{\lambda}$ and $\bm{n}_{0}$ into blocks as
$\bm{\lambda}=\begin{pmatrix}\bm{\lambda}_{1}&\bm{0}\\ \bm{0}&\bm{0}\\ \end{pmatrix},\;\bm{n}_{0}=\begin{pmatrix}\bm{n}_{1}\\ \bm{0}\end{pmatrix},$ (A.6)
where $\bm{\lambda}_{1}$ is diagonal and invertible, and the number of zero rows is $\ell$. Correspondingly, we divide $\bm{P}^{t}$ into two blocks with $n_{\rm tod}-\ell$ and $\ell$ rows, respectively:
$\bm{P}^{t}=\begin{pmatrix}\bm{P}^{t}_{1}\\ \bm{P}^{t}_{2}\end{pmatrix}\;\longrightarrow\;\bm{P}^{t}\bm{d}=\begin{pmatrix}\bm{P}^{t}_{1}\bm{d}\\ \bm{P}^{t}_{2}\bm{d}\end{pmatrix}=\begin{pmatrix}\bm{d}_{1}\\ \bm{d}_{2}\end{pmatrix}.$ (A.13)
Thus, eq. (A.1) becomes
$\begin{pmatrix}\bm{d}_{1}\\ \bm{d}_{2}\end{pmatrix}=\begin{pmatrix}\bm{P}^{t}_{1}\bm{A}\bm{x}\\ \bm{P}^{t}_{2}\bm{A}\bm{x}\end{pmatrix}+\begin{pmatrix}\bm{\lambda}_{1}\bm{n}_{1}\\ \bm{0}\end{pmatrix}.$ (A.20)
Apparently, this equation should first be solved for the lower part because it effectively contains no noise:
$\bm{d}_{2}=(\bm{P}^{t}_{2}\bm{A})\bm{x}.$ (A.21)
However, $\bm{P}^{t}_{2}\bm{A}$ has $\ell$ rows and $n_{\rm{pix}}$ columns, and $\ell$ is almost always much smaller than $n_{\rm{pix}}$, so the number of unknowns is bigger than the number of conditions. Therefore, the equation cannot be solved directly. Here we first take the singular value decomposition of $\bm{P}^{t}_{2}\bm{A}$ as
$\bm{P}^{t}_{2}\bm{A}=\bm{F}\bm{\Lambda}\bm{G}^{t},$ (A.22)
where $\bm{F}$ is $\ell\times\ell$ and unitary; $\bm{G}$ is $n_{\rm{pix}}\times n_{\rm{pix}}$ and unitary; and $\bm{\Lambda}$ has $\ell$ rows and $n_{\rm{pix}}$ columns. However, most elements of $\bm{\Lambda}$ are zero, except for an $\ell\times\ell$ square block $\bm{\Lambda}_{1}$ on its top-left containing the singular values of $\bm{P}^{t}_{2}\bm{A}$.
Let $\bm{x}_{1}=\bm{G}^{t}\bm{x}$ be a column vector, whose first $\ell$ elements are denoted as $\bm{x}_{1,upper}$, and the rest, $\bm{x}_{1,lower}$. We write $\bm{P}^{t}_{2}\bm{A}\bm{x}$ explicitly as:
$\bm{d}_{2}=\bm{P}^{t}_{2}\bm{A}\bm{x}=\bm{F}\begin{pmatrix}\bm{\Lambda}_{1}&\bm{0}\end{pmatrix}\bm{G}^{t}\bm{x}=\bm{F}\begin{pmatrix}\bm{\Lambda}_{1}&\bm{0}\end{pmatrix}\begin{pmatrix}\bm{G}_{1}^{t}\bm{x}\\ \bm{G}_{2}^{t}\bm{x}\end{pmatrix}=\bm{F}\begin{pmatrix}\bm{\Lambda}_{1}&\bm{0}\end{pmatrix}\begin{pmatrix}\bm{x}_{1,\rm{upper}}\\ \bm{x}_{1,\rm{lower}}\end{pmatrix}=\bm{F}\bm{\Lambda}_{1}\bm{x}_{1,\rm{upper}}.$ (A.30)
Thus, eq. (A.21) becomes
$\bm{d}_{2}=\bm{F}\bm{\Lambda}_{1}\bm{x}_{1,\rm{upper}}\;\;\longrightarrow\;\;\bm{x}_{1,\rm{upper}}=\bm{\Lambda}_{1}^{-1}\bm{F}^{t}\bm{d}_{2},$ (A.31)
which is a precise solution (noiseless). Now we can return to eq. (A.20) to deal with its upper half: $\bm{d}_{1}=\bm{P}_{1}^{t}\bm{A}\bm{x}+\bm{\lambda}_{1}\bm{n}_{1}$. Substitute $\bm{x}=\bm{G}\bm{x}_{1}$, we get
$\bm{d}_{1}=(\bm{P}_{1}^{t}\bm{A}\bm{G})\bm{x}_{1}+\bm{\lambda}_{1}\bm{n}_{1}=\bm{A}_{1}\begin{pmatrix}\bm{x}_{1,\rm{upper}}\\ \bm{x}_{1,\rm{lower}}\end{pmatrix}+\bm{\lambda}_{1}\bm{n}_{1},$ (A.34)
Because $\bm{x}_{1,upper}$ is solved precisely, we divide $\bm{A}_{1}$ into blocks with $\ell$ and $n_{\rm{pix}}-\ell$ columns, respectively:
$\bm{A}_{1}=\begin{pmatrix}\bm{A}_{1,\rm{left}}&\bm{A}_{1,\rm{right}}\end{pmatrix},$ (A.36)
then we have
$\bm{d}_{1}=\begin{pmatrix}\bm{A}_{1,\rm{left}}&\bm{A}_{1,\rm{right}}\end{pmatrix}\begin{pmatrix}\bm{x}_{1,\rm{upper}}\\ \bm{x}_{1,\rm{lower}}\end{pmatrix}+\bm{\lambda}_{1}\bm{n}_{1}=\bm{A}_{1,\rm{left}}\bm{x}_{1,\rm{upper}}+\bm{A}_{1,\rm{right}}\bm{x}_{1,\rm{lower}}+\bm{\lambda}_{1}\bm{n}_{1}.$ (A.40)
Because $\bm{x}_{1,\rm{upper}}$ is already known, we get
$\bm{d}_{1}-\bm{A}_{1,left}\bm{x}_{1,\rm{upper}}=\bm{d}_{1}^{\prime}=\bm{A}_{1,\rm{right}}\bm{x}_{1,\rm{lower}}+\bm{\lambda}_{1}\bm{n}_{1}.$ (A.41)
The noise covariance matrix in the above equation is apparently non-singular. Therefore, the problem of singular covariance matrix is solved, and in the next step, we can start with the above equation that has no covariance matrix singularity. Once $\bm{x}_{1,\rm{lower}}$ is solved, we immediately get $\bm{x}=\bm{G}\bm{x}_{1}$.
A.2 How to safely reduce the linear system’s size
In the above section, we have explained how to safely convert the problem with a singular covariance matrix to one with invertible covariance matrix; therefore, below we will assume the noise covariance matrix is invertible, and the noise is already converted to standard white noise like:
$(\bm{\lambda}^{\times}\bm{P}^{t}\bm{d})$ $=(\bm{\lambda}^{\times}\bm{P}^{t}\bm{A})\bm{x}+\bm{n}_{0}.$ (A.42)
For convenience, we rewrite the above equation and renew the symbols $\bm{A}$ and $\bm{d}$, so as to continue with the equation below:
$\bm{d}$ $=\bm{A}\bm{x}+\bm{n}_{0}.$ (A.43)
To obtain the solution of eq. (A.43), we first write the singular value decomposition (SVD) of $\bm{A}$ as
$\bm{A}=\bm{U}\bm{D}\bm{V}^{t},$ (A.44)
where $\bm{U}$ has the same shape as $\bm{A}$ and $\bm{U}^{t}\bm{U}$ is an $n_{\rm{pix}}\times n_{\rm{pix}}$ identity matrix, $\bm{D}$ is an $n_{\rm{pix}}\times n_{\rm{pix}}$ diagonal matrix containing the singular values of $\bm{A}$, and $\bm{V}$ is $n_{\rm{pix}}\times n_{\rm{pix}}$ and unitary. Substitute the decomposition into eq. (A.43) shows
$\bm{U}^{t}\bm{d}$ $=\bm{D}\bm{V}^{t}\bm{x}+\bm{U}^{t}\bm{n}_{0}.$ (A.45)
This operation targets at reducing the size of the problem because, in the above equation, the sizes of the column vectors $\bm{U}^{t}\bm{d}$, $\bm{V}^{t}\bm{x}$ and $\bm{U}^{t}\bm{n}$ are all $n_{\rm{pix}}$, which is much smaller than $n_{\rm tod}$. Meanwhile, because $\bm{U}$ comes from the singular value decomposition of $\bm{A}$, the above step caused no loss of the sky map information from $\bm{A}\bm{x}$ to $\bm{D}\bm{V}^{t}\bm{x}$.
However, there is another concern: Because $\bm{U}$ is not an unitary matrix (non-square), when we left-multiply $\bm{U^{t}}$ to both sides of the equation, it does change the noise information. In order to explain this issue, we introduce another equivalent convention of the singular value decomposition that shows
$\bm{A}=\bm{U}_{1}\bm{D}_{1}\bm{V}^{t}_{1},$ (A.46)
where $\bm{U}_{1}$ is $n_{\rm tod}\times n_{\rm tod}$ and unitary; $\bm{D}_{1}$ has $n_{\rm tod}$ rows and $n_{\rm{pix}}$ columns, but most of its elements are zero, except for the $n_{\rm{pix}}\times n_{\rm{pix}}$ square block on top of it, which is diagonal and contains the singular values of $\bm{A}$; and $\bm{V}_{1}$ is still $n_{\rm{pix}}\times n_{\rm{pix}}$ and unitary. Compared with eq. (A.44), this form append normalized and linearly independent columns to the right of $\bm{U}$ to make it unitary, and zeros to the bottom of $\bm{D}$ to match the rule of matrix multiplication, but do not change the matrix multiplication. Thus, the two forms of singular value decomposition are equivalent.
The advantage of eq. (A.46) is that $\bm{U}_{1}$ is unitary; thus it is always safe to left-multiply $\bm{U}_{1}^{t}$ to both sides of eq. (A.43) to give
$\bm{U}_{1}^{t}\bm{d}=\bm{D}_{1}\bm{V}_{1}^{t}\bm{x}+\bm{U}_{1}^{t}\bm{n}_{0}.$ (A.47)
The above equation contains much more rows than eq. (A.45); however, it is easy to observe the following facts:
  1. 1.
    The top $n_{\rm{pix}}$ lines are identical between eq. (A.45 & A.47).
  2. 2.
    The rest lines in eq. (A.47) contain only noise and no information from the sky map.
  3. 3.
    The covariance matrix of $\bm{U}_{1}^{t}\bm{n}_{0}$ is $\langle(\bm{U}_{1}^{t}\bm{n}_{0})(\bm{U}_{1}^{t}\bm{n}_{0})^{t}\rangle=\bm{I}$, i.e., $\bm{U}_{1}^{t}\bm{n}_{0}$ is another realization of standard white noise.
Items 12 ensure that the sky map information should be solved from the first $n_{\rm{pix}}$ equations, which is exactly eq. (A.45). Item 3 ensures that the noise terms in the rest $n_{\rm tod}-n_{\rm{pix}}$ equations are uncorrelated with the noise terms in the first $n_{\rm{pix}}$ equations; thus, they cannot be used to cancel any noise component in the first $n_{\rm pix}$ rows, i.e., they cannot help to improve the signal-to-noise ratio (SNR). Therefore, the optimal solution given by eq. (A.45, A.47) are identical.
A.3 How to deal with the main matrix’s singularity
Because the main matrix $\bm{A}$ is non-square, its singularity is better represented by the singularity of $\bm{A}^{t}\bm{A}=\bm{V}\bm{D}^{2}\bm{V}^{t}$. When the latter square matrix is singular, there could be zero diagonal elements in $\bm{D}$, like:
$\bm{D}=\begin{pmatrix}\bm{D}_{1}&\bm{0}\\ \bm{0}&\bm{0}\end{pmatrix},$ (A.50)
where $\bm{D}_{1}$ is an invertible diagonal matrix, the number of zero rows is $k$, and all $\bm{0}$s are zero matrices. We then divide the column vectors in eq. (A.45) as follows:
$\bm{U}^{t}\bm{d}=\begin{pmatrix}\bm{d}_{1}\\ \bm{*}\end{pmatrix},\bm{V}^{t}\bm{x}=\begin{pmatrix}\bm{x}_{1}\\ \bm{*}\end{pmatrix},\bm{U}^{t}\bm{n}_{0}=\begin{pmatrix}\bm{n}_{1}\\ \bm{*}\end{pmatrix},$ (A.57)
where “$*\bm{*}$” corresponds to the zero rows of $\bm{D}$, and eq. (A.45) becomes
$\begin{pmatrix}\bm{d}_{1}\\ \bm{*}\end{pmatrix}=\begin{pmatrix}\bm{D}_{1}&\bm{0}\\ \bm{0}&\bm{0}\end{pmatrix}\begin{pmatrix}\bm{x}_{1}\\ \bm{*}\end{pmatrix}+\begin{pmatrix}\bm{n}_{1}\\ \bm{*}\end{pmatrix}=\begin{pmatrix}\bm{D}_{1}\bm{x}_{1}\\ \bm{0}\end{pmatrix}+\begin{pmatrix}\bm{n}_{1}\\ \bm{*}\end{pmatrix}.$ (A.70)
As explained in section A.2, because $\bm{U}^{t}\bm{n}_{0}$ is another realization of standard white noise, the “$*\bm{*}$” rows can be safely removed from the solution; thus the above equation is equivalent to
$\bm{d}_{1}=\bm{D}_{1}\bm{x}_{1}+\bm{n}_{1}.$ (A.71)
Because $\bm{D}_{1}$ is square, diagonal and invertible. It is easy to prove that the optimal solution of the above equation is simply
$\widetilde{\bm{x}}_{1}=\bm{D}_{1}^{-1}\bm{d}_{1}.$ (A.72)
However, note that this solution is in a specific space associated with the unitary matrix $\bm{V}$, and we eventually want to solve the sky map $\bm{x}$ in the ordinary pixel domain.
From eq. (A.72) we see that
$\widetilde{\bm{x}}_{1}=\bm{D}_{1}^{-1}\bm{d}_{1}=\bm{x}_{1}+\bm{D}_{1}^{-1}\bm{n}_{1}.$ (A.73)
With a similar block scheme as eq. (A.57), we define
$\bm{V}^{t}=\begin{pmatrix}\bm{V}_{1}^{t}\\ \bm{*}\end{pmatrix},\bm{U}^{t}=\begin{pmatrix}\bm{U}_{1}^{t}\\ \bm{*}\end{pmatrix},$ (A.78)
where $\bm{V}_{1}^{t}$ is a matrix consist of the first $n_{\rm{pix}}-k$ rows of $\bm{V}^{t}$, $\bm{U}_{1}^{t}$ is a matrix consist of the first $n_{\rm{pix}}-k$ rows of $\bm{U}^{t}$, and $k$ is the number of zero rows in eq. (A.50); so we have
$\bm{x}_{1}=\bm{V}_{1}^{t}\bm{x},\;\bm{n}_{1}=\bm{U}_{1}^{t}\bm{n},$ (A.79)
and we get the following linear equations that consist of $n-k$ rows to solve $\bm{x}$ from.
$\widetilde{\bm{x}}_{1}=\bm{V}_{1}^{t}\bm{x}+\bm{D}_{1}^{-1}\bm{U}_{1}^{t}\bm{n}.$ (A.80)
The goal is to estimate $\bm{x}$ from the above equation, and the main difficulty is that $\bm{V}_{1}^{t}$ has $n_{\rm{pix}}-k$ rows but $n_{\rm{pix}}$ columns, which is insufficient to solve $\bm{x}$ fully. However, because $\widetilde{\bm{x}}_{1}$, $\bm{V}_{1}^{t}\bm{x}$ and $\bm{D}_{1}^{-1}\bm{U}_{1}^{t}\bm{n}$ are all column vectors of size $n_{\rm{pix}}-k$, there is no problem to add $k$ zero elements to the bottom of each of them to expand their sizes from $n_{\rm{pix}}-k$ to $n_{\rm{pix}}$. Then we can rewrite the equation as
$\widetilde{\bm{x}}_{2}=\bm{V}^{t}\bm{x}+\bm{n}_{2},$ (A.81)
where
$\widetilde{\bm{x}}_{2}=\begin{pmatrix}\widetilde{\bm{x}}_{1}\\ \bm{0}\end{pmatrix},\bm{n}_{2}=\begin{pmatrix}\bm{D}_{1}^{-1}\bm{U}_{1}^{t}\bm{n}\\ \bm{0}\end{pmatrix},$ (A.86)
and we require a special constraint on $\bm{x}$ that its inner products with the last $k$ rows of $\bm{V}^{t}$ (which is an $n_{\rm{pix}}\times n_{\rm{pix}}$ unitary matrix) are zero, which means the corresponding modes are supposed to be missing in $\bm{x}$.
Because $\bm{V}^{t}$ is $n_{\rm{pix}}\times n_{\rm{pix}}$ and unitary, it is easy to prove that the optimal solution of eq. (A.81) is simply
$\widetilde{\bm{x}}=\bm{V}\widetilde{\bm{x}}_{2}=\bm{V}\begin{pmatrix}\bm{D}_{1}^{-1}\bm{d}_{1}\\ \bm{0}\end{pmatrix},$ (A.89)
which is equivalent to the following:
$\widetilde{\bm{x}}=\bm{V}\begin{pmatrix}\bm{D}_{1}^{-1}&\bm{0}\\ \bm{0}&\bm{0}\end{pmatrix}\bm{U}^{t}\bm{d}$ (A.92)
In fact, eq. (A.81A.92) are mainly used to derive the solution. Once we obtain eq. (A.92), there is no problem to proceed to the final form of the solution that is free from zero rows or columns. Because both $\bm{U}$ and $\bm{V}$ can be divide in columns as
$\bm{U}=\begin{pmatrix}\bm{U}_{1}&\bm{*}\end{pmatrix},\,\bm{V}=\begin{pmatrix}\bm{V}_{1}&\bm{*}\end{pmatrix};$ (A.95)
thus,
$\widetilde{\bm{x}}=\begin{pmatrix}\bm{V}_{1}&\bm{*}\end{pmatrix}\begin{pmatrix}\bm{D}_{1}^{-1}&\bm{0}\\ \bm{0}&\bm{0}\end{pmatrix}\begin{pmatrix}\bm{U}_{1}^{t}\\ \bm{*}\end{pmatrix}\bm{d}=\bm{V}_{1}\bm{D}_{1}^{-1}\bm{U}_{1}^{t}\bm{d}.$ (A.101)
The above equation give the optimal solution of map-making, provided that the covariance matrix singularity has already been solved by the general approach in section A.1.
A.4 Features of the optimal solution
Provided we have no prior knowledge of the missing modes of the input sky map [^2] The term “no prior knowledge” means the missing modes are regarded as lost forever, and there is no additional knowledge or information that can help to recover them., the optimal solution is the unique solution that satisfy the three important features listed below:
  1. 1.
    Lossless: $\widetilde{\bm{x}}$ is lossless except for the modes that are lost before any solution (prior loss).
  2. 2.
    Minimum error: The TOD-domain estimation error $\bm{\delta}=\bm{d}-\bm{A}\widetilde{\bm{x}}$ is minimized.
  3. 3.
    Minimum length: Of all solutions that satisfy items 12, $\widetilde{\bm{x}}$ (ignore noise) has the lowest pixel domain length (or power).
Therefore, $\widetilde{\bm{x}}$ is the best blind solution (BBL) of the problem, where “blind” explicitly refers to the pre-condition that we have no prior knowledge of the missing modes of the sky map.
In a less complicated scenario where singular noise covariance matrices do not need to be fully addressed, a solution analogous to the above was described by [26]. Nonetheless, their methodology relied on the employment of an eigenvalue decomposition, which may not be applicable to non-square matrices and could potentially falter when dealing with certain square matrices, such as those derived from the Jordan normal form. Moreover, the eigen-problem of a real-valued system commonly necessitates resolution within the complex space. Meanwhile, it is also possible to deal with the singularity problem by means of limits, like the one shown in [38], which is often implemented by the use of an additional small matrix; however, this will inevitably propagate the side effect to all modes, which will not happen in the above general solution. Even if for the simplified solution in section A.5.2, the side effect will be strictly limited within the missing modes.
A.5 Possible simplifications
In this section, we discuss possible simplifications to the general optimal solution.
A.5.1 With tightly associated singularities
As shown in section A.1, it is quite complicated to deal with a singular covariance matrix in general. However, in some cases, the singularity of the covariance matrix comes from the posterior processing of TOD, i.e., the singularities of matrix $\bm{A}$ and $\bm{C}$ are tightly associated, which can make the problem easier. For example, assume both $\bm{A}^{t}\bm{A}$ and $\bm{C}$ are non-singular at the beginning, but the TOD is later processed as
$\bm{d}\rightarrow\bm{M}\bm{d}=\bm{M}\bm{A}\bm{x}+\bm{M}\bm{n},$ (A.102)
where $\bm{M}$ is singular. One can prove that in this case, the lower parts of eq. (A.20) are all zero and can be safely ignored to leaves only the upper half
$\bm{P}^{t}_{1}\bm{d}=\bm{P}^{t}_{1}\bm{A}\bm{x}+\bm{\lambda}_{1}\bm{n}_{1}\;\;\Longrightarrow\;\;\bm{d}_{1}=\bm{A}_{1}\bm{x}+\bm{\lambda}_{1}\bm{n}_{1},$ (A.103)
whose noise covariance matrix is non-singular.
In the above mentioned special case, one can prove that the optimal solution is similar to eq.(2.2), just replace each matrix inversion with the pseudo inverse:
$\widetilde{\bm{x}}=(\bm{A}^{t}\bm{C}^{+}\bm{A})^{+}\bm{A}^{t}\bm{C}^{+}\bm{d}.$ (A.104)
It should be noted that even if the singularity arises solely from filtering, the matrix $\bm{A}$ and $\bm{C}$ are not always associated. This is especially true if the filtering involves a subtractive operation, such as
$\bm{d}\rightarrow\bm{d}^{\prime}=\bm{A}\bm{x}+\bm{n}-\sum_{i}\bm{T}_{i},$ (A.105)
Here, $\bm{T}_{i}$ refers to the external templates that are used to eliminate or mitigate contaminations. Estimating the impact of subtraction on the noise and signal covariance matrices is challenging because their cross-correlation is often unpredictable, and especially, for a blind estimate, the signal is assumed to be unknown. However, it is necessary to consider the possibility that the noise and signal covariance matrices may be affected differently.
A.5.2 The application of modified pseudo inverse
Further more, one can also use the modified pseudo inverse to deal with a singular noise covariance matrix, which can also significantly simplify the processes but does not require the singularity association between $\bm{A}$ and $\bm{C}$.
The modified pseudo inverse is defined as follows: Let $\lambda_{i}$ be the diagonal elements of a diagonal matrix $\bm{\lambda}$, the modified pseudo inverse of $\bm{\lambda}$ is another diagonal matrix $\bm{\lambda}^{\times}$ of the same shape, whose diagonal elements $\lambda_{i}^{\times}$ are given by
$\lambda_{i}^{\times}=\left\{\begin{matrix}\lambda_{i}^{-1}&(\lambda_{i}\neq 0)\\ 1&(\lambda_{i}=0)\\ \end{matrix}\right.$ (A.108)
Apparently, if we change the above definition to $\lambda_{i}^{\times}=0$ when $\lambda_{i}=0$, then the definition returns to the pseudo inverse, usually noted as $\bm{\lambda}^{+}$. For a general non-diagonal matrix $\bm{M}$, the pseudo and modified pseudo inverses are defined based on its singular-value decomposition: Assume $\bm{M}$’s singular value decomposition is $\bm{M}=\bm{V}\bm{\lambda}\bm{U}^{t}$, and then its pseudo and modified pseudo inverses are:
$\bm{M}^{+}$ $=\bm{V}\bm{\lambda}^{+}\bm{U}^{t}$ (A.109)
$\bm{M}^{\times}$ $=\bm{V}\bm{\lambda}^{\times}\bm{U}^{t},$
respectively. The main difference between $\bm{\lambda}^{\times}$ and $\bm{\lambda}^{+}$ is that the modified pseudo inverse $\bm{\lambda}^{\times}$ is full ranked (invertible); thus, it is always safe to left-multiply $\bm{\lambda}^{\times}$ to both sides of eq. (A.1) to get
$\bm{\lambda}^{\times}\bm{P}^{t}\bm{d}$ $=\bm{\lambda}^{\times}\bm{P}^{t}\bm{A}\bm{x}+\begin{pmatrix}\bm{n}_{1}\\ \bm{0}\end{pmatrix}.$ (A.112)
Then we replace the noise term in the above equation with $\bm{n}_{0}$ for approximation, and get
$\bm{\lambda}^{\times}\bm{P}^{t}\bm{d}$ $\approx\bm{\lambda}^{\times}\bm{P}^{t}\bm{A}\bm{x}+\bm{n}_{0},$ (A.113)
which eliminates the covariance matrix singularity at the price of virtually injecting the following difference into the linear system:
$\begin{pmatrix}\bm{n}_{1}\\ \bm{0}\end{pmatrix}-\bm{n}_{0}=\begin{pmatrix}\bm{0}\\ \bm{\delta}\end{pmatrix},$ (A.118)
where $\bm{\delta}$ is a standard white noise with $\ell$ rows. Therefore, the injection is strictly limited within the singular part, and the energy of the virtually injected noise is relatively $\ell/(n_{\rm tod}-\ell)$. Because $\ell\ll n_{\rm tod}$, the noise injection is negligible in the vast majority of cases, and will not significantly deteriorate the solution. Especially, because the noise injection is only “virtually” (no actual injection), the side effect is further reduced, making the corresponding solution near optimal.
When the modified pseudo inverse is adopted, the near-optimal solution can be shown as
$\widetilde{\bm{x}}=(\bm{A}^{t}\bm{C}^{\times}\bm{A})^{+}\bm{A}^{t}\bm{C}^{\times}\bm{d},$ (A.119)
which is similar to eq. (A.104), only that the inversion of the covariance matrix is given by the modified pseudo inverse rather than the standard pseudo inverse.
A.6 Summary of the optimal and recommended solutions
As described above, the true optimal solution should be given by the following approach:
  1. 1.
    First appropriately deal with the covariance matrix singularity as shown in section A.1.
  2. 2.
    Then follow sections A.2A.3 to solve the rest of the problem.
Because the true optimal solution is quite complex, two simplified solutions are given above to make life easier. The solution with pseudo inverse (section A.5.1) requires a special precondition to be valid, and will also cause more problem regarding the scan strategy (to be discussed in section 3.1). The near-optimal solution with modified pseudo inverse (section A.5.2) required no precondition, has strictly limited side effect, and is more robust. Therefore, the near-optimal solution in eq.(A.119) is recommended.
Compared with the well known minimum variance solution of the simplest case in eq. (2.2), the recommended solution in eq. (A.119) inverts the noise covariance matrix $\bm{C}$ by the modified pseudo inverse, and other matrices by the standard pseudo inverse. As discussed in section A.5.2, the side effect of eq. (A.119) is strictly limited in the singular part. Thus the difference to the true optimal solution is almost always negligible.
Another advantage of eq. (A.119) is that it is suitable for fast computation, for several reasons:
  1. 1.
    There is no need to compute the SVD of the huge matrix $\bm{A}$.
  2. 2.
    The inversion of $\bm{A}^{t}\bm{C}^{\times}\bm{A}$ is for an $n_{\rm{pix}}\times n_{\rm{pix}}$ matrix, not an $n_{\rm tod}\times n_{\rm tod}$ matrix, which is much easier.
  3. 3.
    Both $\bm{A}^{t}\bm{C}^{\times}\bm{A}$ and the covariance matrix are symmetric, which is easier to handle.
  4. 4.
    The inversion of the noise covariance matrix $\bm{C}$ can be done by e.g., the Fast Fourier transform, which is much faster than matrix multiplication.
  5. 5.
    Because eq. (A.119) prevents the information loss on the CMB signal as much as possible, it does not cause any unrecoverable E-to-B leakage by itself, which is an important advantage for detecting the primordial gravitational waves via CMB.
We also point out an interesting fact about the near-optimal solution: According to section A.1 and especially eq. (A.31), if $\ell$ (the number of zero eigenvalue of $\bm{C}$) is sufficiently big, then eq. (A.31) may contain enough noiseless rows to solve $\bm{x}$ completely and precisely, making all rest of the equations useless. The true optimal solution can properly deal with this case, hence it is definitely better than other solutions. However, in most cases, $\ell$ must be very small, which makes the difference between the optimal and near-optimal solution negligible. This is the precondition for eq. (A.119) to be useful.
A.7 A quick reference
This section is only a quick reference for busy readers, without any new content.
For the problem of solving $\bm{x}$ from $\bm{d}=\bm{A}\bm{x}+\bm{n}$, the optimal solution is given as follows: First follow the general approach in section A.1 to deal with the singularity of the covariance matrix and convert the noise to standard white noise to obtain a renewed main equation:
$\bm{d}=\bm{A}\bm{x}+\bm{n}_{0},$ (A.120)
which still allows a singular $\bm{A}^{t}\bm{A}$. The optimal solution of the above renewed equation is precisely:
$\widetilde{\bm{x}}=(\bm{A}^{t}\bm{A})^{+}\bm{A}^{t}\bm{d}.$ (A.121)
Practically, if the singularity of the covariance matrix is handled by the modified pseudo inverse, then we get the near-optimal solution directly from the original linear system $\bm{d}=\bm{A}\bm{x}+\bm{n}$ as
$\widetilde{\bm{x}}\approx(\bm{A}^{t}\bm{C}^{\times}\bm{A})^{+}\bm{A}^{t}\bm{C}^{\times}\bm{d}.$ (A.122)
Although the solution in eq. (A.122) may be slightly sub-optimal, it is recommended because it involves a considerably easier computational procedure than the optimal option. Additionally, the divergence from the optimal solution is strictly limited to the singular components, which us usually insignificant.
Appendix B Efficient computation of the noise covariance matrix’s modified pseudo inverse
To calculate the sub-optimal solution outlined in equation (A.122), it is necessary to determine the modified pseudo-inverse of matrix $\bm{C}$. However, it should be noted that $\bm{C}$ is a matrix in the time domain, which may have a significantly larger size compared to a pixel domain matrix. Consequently, it becomes imperative to explore efficient methods for computing $\bm{C}^{\times}$.
Usually, $\bm{C}$ can be regarded as diagonal in the Fourier domain (at least approximately), so its time-domain pseudo inverse is simply the following:
$\bm{C}=\bm{W}\bm{\lambda}^{2}\bm{W}^{H}\longrightarrow\bm{C}^{\times}=\bm{W}(\bm{\lambda}^{2})^{\times}\bm{W}^{H},$ (B.1)
where $\bm{W}$ is the Fourier transform matrix. However, when the noise is filtered like $\bm{n}\longrightarrow\bm{M}\bm{n}$, the above equation can no longer be used to compute the modified pseudo inverse of $\bm{C}$. In this case, we first consider the singular value decomposition of the filtering matrix as $\bm{M}=\bm{U}\bm{\Lambda}\bm{V}^{H}$, and convert the time domain noise to white noise as shown in eq. (A.1), $\bm{n}=\bm{P}\bm{\lambda}\bm{n}_{0}$, to obtain
$\bm{d}=\bm{A}\bm{x}+\bm{M}\bm{n}=\bm{A}\bm{x}+\bm{U}\bm{\Lambda}\bm{V}^{H}\bm{P}\bm{\lambda}\bm{n}_{0}.$ (B.2)
Left-multiply $\bm{U}^{H}$ to both sides:
$\bm{U}^{H}\bm{d}=\bm{U}^{H}\bm{A}\bm{x}+\bm{\Lambda}\bm{V}^{H}\bm{P}\bm{\lambda}\bm{n}_{0},$ (B.3)
then with the idea of modified pseudo inverse, left-multiply $\bm{\Lambda}^{\times}$ to both sides:
$\bm{\Lambda}^{\times}\bm{U}^{H}\bm{d}=\bm{\Lambda}^{\times}\bm{U}^{H}\bm{A}\bm{x}+\bm{\Lambda}^{\times}\bm{\Lambda}\bm{V}^{H}\bm{P}\bm{\lambda}\bm{n}_{0},$ (B.4)
and assume $\bm{\Lambda}^{\times}\bm{\Lambda}\approx\bm{I}$. As explained in section A.5.2, this means to assume some additional noise in the singular part, whose side effect is strictly limited and usually insignificant. With the approximation, we further get
$\bm{\lambda}^{\times}\bm{P}^{H}\bm{V}\bm{\Lambda}^{\times}\bm{U}^{H}\bm{d}\approx\bm{\lambda}^{\times}\bm{P}^{H}\bm{V}\bm{\Lambda}^{\times}\bm{U}^{H}\bm{A}\bm{x}+\bm{n}_{0},$ (B.5)
whose solution is
$\widetilde{\bm{x}}=(\bm{A}_{1}^{H}\bm{A}_{1})^{+}\bm{A}_{1}^{H}\bm{d}_{1},$ (B.6)
which does not require to compute $\bm{C}^{\times}$, and
$\bm{A}_{1}$ $=\bm{\lambda}^{\times}\left(\bm{P}^{H}\bm{V}\bm{\Lambda}^{\times}\bm{U}^{H}\right)\bm{A}=\bm{\lambda}^{\times}(\bm{P}^{H}\bm{M}^{\times})\bm{A}$ (B.7)
$\bm{d}_{1}$ $=\bm{\lambda}^{\times}\left(\bm{P}^{H}\bm{V}\bm{\Lambda}^{\times}\bm{U}^{H}\right)\bm{d}=\bm{\lambda}^{\times}(\bm{P}^{H}\bm{M}^{\times})\bm{d}.$
As long as the filtering scheme does not change, $\bm{M}$ is a constant matrix, so do $\bm{M}^{\times}$ and $\bm{P}^{H}\bm{M}^{\times}$. Therefore, with eqs. (B.6B.7), the time domain eigen-problem is converted to matrix multiplication, which can be computed efficiently once the noise spectrum $\bm{\lambda}$ is obtained.
Appendix C More discussion about the singularity of the main matrix
It is not always easy to determine the singularity of matrix $\bm{A}^{t}\bm{C}^{+}\bm{A}$ when $\bm{C}^{+}$ is a singular matrix. To determine the rank of matrix $\bm{A}^{t}\bm{C}^{+}\bm{A}$, we can find the dimension of the null-space of it using the rank-nullity theorem.
If $\bm{v}\in{\rm Null}(\bm{A}^{t}\bm{C}^{+}\bm{A})$, then it implies that
$\bm{v}^{t}\bm{A}^{t}\bm{C}^{+}\bm{A}\bm{v}=0.$ (C.1)
Assume that $\bm{A}$ has full column-rank, which means for any $\bm{v}\neq\bm{0}$, $\bm{Av}\neq\bm{0}$. Thus $\bm{Av}\in{\rm Null}(\bm{C}^{+})$. Due to the fact that $\bm{C}^{+}$ and $\bm{C}$ have the same null-space because $\bm{C}$ is symmetric, it is evident that
$\bm{v}^{t}\bm{A}^{t}\bm{C}\bm{A}\bm{v}=0.$ (C.2)
Therefore, we can conclude that ${\rm Null}(\bm{A}^{t}\bm{C}^{+}\bm{A})\subseteq{\rm Null}(\bm{A}^{t}\bm{C}\bm{A})$. On the contrary, assuming that $\bm{v}\in{\rm Null}(\bm{A}^{t}\bm{C}\bm{A})$, we can also conclude that ${\rm Null}(\bm{A}^{t}\bm{C}\bm{A})\subseteq{\rm Null}(\bm{A}^{t}\bm{C}^{+}\bm{A})$. These two conclusions indicate that $\bm{A}^{t}\bm{C}^{+}\bm{A}$ and $\bm{A}^{t}\bm{C}\bm{A}$ have the same null-space.
Usually, the singularity of matrix $\bm{C}$ is caused by filtering. When applying a filter to TOD, the covariance matrix will be modified, and this modification can be represented as $\bm{C}=\bm{M}\tilde{\bm{C}}\bm{M}^{t}$, where $\tilde{\bm{C}}$ is a positive-definite matrix, and $\bm{M}$ represents the filter. Again, if $\bm{v}\in{\rm Null}(\bm{A}^{t}\bm{C}\bm{A})$, we have
$\bm{v}^{t}\bm{A}^{t}\bm{M}\tilde{\bm{C}}\bm{M}^{t}\bm{A}\bm{v}=0.$ (C.3)
Since $\tilde{C}$ is positive-definite, we can state that the equation holds if and only if $\bm{M}^{t}\bm{Av}=\bm{0}$, which also means that $\bm{A}^{t}\bm{C}\bm{A}$ and $\bm{M}^{t}\bm{A}$ have the same null-space. Thus $\bm{A}^{t}\bm{C}^{+}\bm{A}$ and $\bm{M}^{t}\bm{A}$ have the same null-space.
Commonly used filters, such as high-pass, polynomial, and Wiener filters, are symmetric, meaning that $\bm{M}=\bm{M}^{t}$. We can write SVD of $\bm{M}$ and $\bm{A}$ as
$\bm{M}$ $=\bm{U}\bm{\lambda}\bm{V}^{t}$ $\bm{A}$ $=\bm{P}\bm{\delta}\bm{Q}^{t}.$ (C.4)
Thus $\bm{MA}=\bm{U\lambda V}^{t}\bm{P\delta Q}^{t}$. Let $\bm{G}$ be represented as $\bm{V}^{t}\bm{P}$. It is evident that $\bm{G}$ is an $n_{\rm tod}$-by-$n_{\rm tod}$ unitary matrix. Divide $\bm{\lambda}$, $\bm{G}$ and $\bm{\delta}$ into block as
$\bm{\lambda}$ $=\begin{pmatrix}\bm{\lambda}_{1}&\bm{0}\\ \bm{0}&\bm{0}\end{pmatrix}$ $\bm{G}$ $=\begin{pmatrix}\bm{G}_{1}&\bm{G}_{2}\\ \bm{G}_{3}&\bm{G}_{4}\end{pmatrix}$ $\bm{\delta}$ $=\begin{pmatrix}\bm{\delta}_{1}\\ \bm{0}\end{pmatrix},$ (C.11)
then the only singular part in $\bm{MA}$ can be denote as
$\bm{\lambda G\delta}=\begin{pmatrix}\bm{\lambda}_{1}\bm{G}_{1}\bm{\delta}_{1}&\bm{0}\\ \bm{0}&\bm{0}\end{pmatrix},$ (C.14)
which means the final rank is determined by the rank of the upper left corner of $\bm{G}$. The upper left part of $\bm{G}$ is formed by multiplying the eigenvector matrix of $\bm{M}$ with the column singular vectors of $\bm{A}$, which is equivalent to performing a Fourier transform on the column singular vectors of $\bm{A}$. Therefore, the final rank depends on the Fourier space structure of the column singular vectors of $\bm{A}$. If their main differences in Fourier space happen to be concentrated in the zeroed-out lower-left part, singularity will occur. For random scanning, the Fourier space structure of the column singular vectors of $\bm{A}$ should be randomly distributed among different components, making it unlikely to be concentrated in the lower-left part. Hence, singularity is hard to appear. However, there is one exception when $\bm{v}$ is proportional to $\bm{1}_{n_{\rm pix}}$. In this case, $\bm{Av}$ is proportional to $\bm{1}_{n_{\rm tod}}$, and if the filter $\bm{M}$ removes the mean value, then $\bm{MAv}=0$.
In summary, if the filter matrix $\bm{M}$ is capable of filtering out the monopole component, the rank of $\bm{A}^{t}\bm{C}^{+}\bm{A}$ is no more than $n_{\rm pix}-1$. However, if the filter matrix $\bm{M}$ does not remove the monopole component, the rank can be $n_{\rm pix}$ in the best case.
Appendix D The code validation
All three solutions $\bm{S}_{+}$, $\bm{S}_{\times}$, and $\bm{S}_{\rm c}$ are unbiased, with the exception of the lost modes. In computing $\bm{C}^{+}$, we set a threshold and ignore the small singular values in order to improve the numerical stability. Assuming that the suppression effect is independent at each $\ell$, the final power spectra can be corrected by running several simulations to get the suppression factor. Then we can evaluate the effectiveness of the three solutions $\bm{S}_{\times},\bm{S}_{+}$, and $\bm{S}_{\rm c}$ by comparing their uncertainties because they are all unbiased for the remaining modes.
In order to validate the properties of $\bm{S}_{+},\bm{S}_{\times}$ and $\bm{S}_{\rm c}$, the magnitudes of the CMB signal and total noise are modulated as $\bm{d}=f_{s}\bm{d}_{\rm cmb}+f_{n}\bm{n}$. In the context of a solution that enables complete reinstatement of the CMB signal (lossless), the solution error is determined only by the noise. Consequently, upon fixing the noise amplitude $f_{n}$ and varying only the CMB amplitude $f_{s}$ during the simulation, the ensuing errors in the $\bm{S}_{\rm c}$ and $S_{\rm\times}$ solutions should not change with $f_{s}$. This holds true regardless of the singularity of the noise covariance matrix $\bm{C}$. However, there should arises a difference in the case of $\bm{S}_{\rm+}$: If the noise covariance matrix $\bm{C}$ is indeed singular, the error in $\bm{S}_{\rm+}$ will exhibit a change with $f_{s}$, because under these specific conditions, $\bm{S}_{\rm+}$ leads to a loss in the CMB signal, and the residual RMS will hence scale as $\sqrt{a^{2}f_{s}^{2}+b^{2}f_{n}^{2}}$. All these expectations are confirmed by the data depicted in Figure 6.
Figure 6: The pixel domain RMS (Root Mean Square) of the residual (output-input) as a function of $f_{s}$, with $f_{n}$ fixed at 1 for a non-singular covariance matrix (left), and a singular covariance matrix (right).
Alternatively, if one maintains a constant CMB amplitude and varies the noise amplitude $f_{n}$ during simulation, the errors in $\bm{S}_{\rm c}$ and $S_{\rm\times}$ should reveal a direct proportionality to $f_{n}$ because there is no CMB error in them (lossless), which applies regardless of the singularity of the noise covariance matrix $\bm{C}$. In contrast, when the noise covariance matrix $\bm{C}$ is singular, the error in $\bm{S}_{\rm+}$ should not exhibit a linear relationship with $f_{n}$, because it contains contributions both from the lost CMB signal and the residual noise. This phenomenon is also confirmed by the data presented in Figure 7. Furthermore, both Figure 67 show that $S_{\rm\times}$ has a significantly lower residual noise than $\bm{S}_{\rm c}$, thereby corroborating its superior efficacy in red noise suppression.
Figure 7: Similar to figure 6 but change $f_{n}$ instead.
The results in Figure 67 are well consistent with expectations, which confirms the validity of our code implementation and ensures reliable results in the tests that follows.